OpenStudy (anonymous):

Find the inverse of f(x)=2e^-x

OpenStudy (saifoo.khan):

\[\Large f^{-1}(x) = -\log(\frac{x}{2})\]

OpenStudy (dumbcow):

it could also be written as \[f^{-1}(x) = \ln 2 - \ln x\]